Two adjacent sides of a parallelogram meet at an angle of 450 and have lengths of 5 and 11 feet respectively. Find the length of the diagonals.


1
Expert's answer
2022-06-06T16:49:39-0400



We can use the Law of Cosines in the triangle ABD:

"BD^2=AB^2+AD^2-2AB\cdot AD cos\angle A,"

"BD^2=5^2+11^2-2\cdot 5\cdot 11 \cdot cos45\degree="

"=25+121-2\cdot 55\cdot \frac{\sqrt{2}}{2}=146-55\sqrt{2},"

"BD=\sqrt{146-55\sqrt{2}}."


Let's find the angle B:

"\angle B=180\degree - \angle A=180\degree -45\degree =135\degree."




And now we can use the Law of Cosines in the triangle ABC:

"AC^2=AB^2+BC^2-2AB\cdot BC cos\angle A,"

"BD^2=5^2+11^2-2\cdot 5\cdot 11 \cdot cos135\degree="

"=25+121-2\cdot 55\cdot (-\frac{\sqrt{2}}{2})=146+55\sqrt{2},"

"AC=\sqrt{146+55\sqrt{2}}."


Answer: "\sqrt{146-55\sqrt{2}}," "\sqrt{146+55\sqrt{2}}."